OpenStudy (anonymous):

Find the slope of the line passing through the points (-3, -8) and (4, 6). (If the answer is a fraction, use the “/” on your keyboard.)

OpenStudy (calculusxy):

\[Slope = \frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (calculusxy):

You can use either one of your coordinates and the x- and y-values to substitute into the variables in the equation.

OpenStudy (calculusxy):

I will use 6 as \(y_2\) and -8 as \(y_1\) and 4 can be \(x_2\) and -3 can be \(x_1\). \[\frac{ 6 - (-8) }{ 4 - (-3) } = \frac{ 14 }{ 7 } = ? \]

OpenStudy (calculusxy):

Do you see how I did that?

OpenStudy (alexandervonhumboldt2):

use formula called slope formula \[\frac{ y_1-y_2 }{ x_1-x_2 }~~or~~\frac{ y_2-y_1 }{ x_2-x_1}\] substitute your points and solve: \[\frac{ -8-6 }{ -3-4 }=\frac{ -14 }{ -7 }=\frac{ 14 }{ 7 }=?\]